OpenStudy (anonymous):

Find the limit of the function by using direct substitution. limit as x approaches zero of x^2-6

OpenStudy (gorv):

x^2-6 as x approch zero 0^2-6=0-6=-6

OpenStudy (anonymous):

thank you! thats what i got just needed to make sure!